OpenStudy (anonymous):

Multiply and simplify 4 plus radical 8 times the quantity 4- radical 8

OpenStudy (anonymous):

\[4 + \sqrt{8} \times (4 - \sqrt{8})\]Is this the problem?

OpenStudy (anonymous):

yes but the first part is in parenthesis too

OpenStudy (anonymous):

So it's \[(4 + \sqrt{8})(4 - \sqrt{8})\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright. Solve what's below and combine like terms. That will be your answer. \[4(4 - \sqrt{8}) + \sqrt{8}(4 - \sqrt{8}) \implies 16 - 4\sqrt{8} + 4\sqrt{8} - 8\implies?\]

OpenStudy (anonymous):

would that be 8+2\[\sqrt{8}\]